OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 6 when x = 4. Enter the constant of variation in lowest terms. y = ?/? x

OpenStudy (danjs):

y = k*x and a point is given (x,y)=(4,6) 6=k*4 k = 6/4 = 3/2 so, y= 3/2 x